OpenStudy (anonymous):

1 Which is the right period, amplitude and phase shift of this trigonometric function y = 6 pie cos ( 3 pie x + 2 )

hartnn (hartnn):

The general equation is \(\large y = Asin(\omega x+\theta)\) where A is the amplitude, Period = T = 2\(\pi\)/\(\omega\) and phase shift = \(\theta\)

hartnn (hartnn):

can u get those parameters now ?

OpenStudy (anonymous):

yes, thank you so much. im stuck with these assignments,

hartnn (hartnn):

welcome, if you don't get correct values, you can ask :)

OpenStudy (anonymous):

heres another :) y = sin x If we shift the function -2 units vertically and 3 units to the right, what would bethe translated function?

hartnn (hartnn):

vertical shift of -2 means, 2 units down, if y was 0 previously, now it will be -2 so, y is replaced by y+2, similarly , x will be replaced by x+3 ok ?

OpenStudy (anonymous):

got it, thank you so much, your an angel :)

hartnn (hartnn):

i am just a helping human :) and welcome ^_^

OpenStudy (anonymous):

hello, help again, i cant get he correct answer. Find the derivative of the function: a) log square( cos(2x) + x cube) and b) log x over 3x square

hartnn (hartnn):

if u tried it, u show your work, i will try to find error.

OpenStudy (anonymous):

i have no scanner here,my papers are all over my bed (:

hartnn (hartnn):

ok, the question is \(\log^2[\cos2x+x^3]\) right ?

OpenStudy (anonymous):

yes